Circular No. 8549

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005cn IN NGC 5061 Further to IAUC 8530, C. Jacques reports the discovery by the Brazil-

ian Supernovae Search Team (including also C. Colesanti, E. Pimentel, and T. Napoleao) of an apparent supernova (mag 14.6) on unfiltered CCD images taken on June 19.99 and 21.00 UT. The new object is located at α = 13^{h}18^{m}00^{s}.46, δ = −26^{o}48′33′′.1 (equinox 2000.0), which is 62′′ west and 103′′ north of the center of NGC 5061. Nothing was visible at this location on a CCD image taken on Apr. 15.12 (limiting mag 19.0) or on a red Digitized Sky Survey image from 1992.

COMETS C/2005 J3–J10 (SOHO) Following are the initial available positions for additional Kreutz sun-

grazing comets found on SOHO website images (continuation to IAUC 8548). K. Battams writes that C/2005 J3, J8, and J10 were faint and diffuse — barely above the background, adding that the peak magnitude for each such object is ≈ 9 ± 0.5; C/2005 J4 was stellar in appearance and also barely above the background. C/2005 J5 had a ‘teardrop’ shape, peaking in brightness at mag ∼ 6.8 at 6.2R_{⊙} (May 9.004 UT). C/2005 J6 was fairly diffuse and reached mag ∼ 6.8 at 6.2R_{⊙} in the C2 coronagraph (May 10.621), but (like C/2005 J5) this object was too close to the pylon for photometry in C3 images; it also displayed a 80′′ tail in the C2 image from May 10.629. C/2005 J7 appeared stellar in both C3 and C2 images, reaching mag ∼ 6.1 at 9.0R_{⊙} (May 11.488, C3). C/2005 J9 was also extremely faint and appeared as a headless tail (230′′ long on May 14.078) at the edge of the field-of-view; as it neared the sun, “it became more of a small, diffuse blob”.

Comet  2005 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2005 J3 May 2.893 2^{h}46.^{m}1 +14^{o}24′ C2  CL 2005-M14
C/2005 J4 5.746 2 56.3 +15 08 C2 MM 2005-M14
C/2005 J5  8.863 3 10.6 +15 46 C3/2 TH 2005-M15
C/2005 J6  10.346 3 17.1 +15 47 C3/2 BZ 2005-M15
C/2005 J7  10.696 3 25.9 +14 36 C3/2 BZ 2005-M15
C/2005 J8  12.724 3 22.4 +16 53 C2  HS 2005-M15
C/2005 J9  14.078 3 28.0 +17 18 C2  JS  2005-M15
C/2005 J10  14.536 3 29.9 +17 19 C2  BZ 2005-M15
2005 June 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green